# The Great Comets of 1843 and 1882 at Their Previous Return to Perihelion in the Twelfth Century: One Spectacular, the Other Dull

### v1

@@ 8. Fragmented Nucleus of the Comet of 1138
This conclusion is supported by Kreutz’s own orbital computations, which for the three separation distances give, respectively, ${}^{\prime\prime}\!$.8, ${}^{\prime\prime}\!$.2, and ${}^{\prime\prime}\!$.6 (Table 8). The distances BA and BC agree with Elkin’s measurements to $\pm$0′′.5 and only CD is off by nearly 2′′, possibly because of a contribution to D from E. Another strong argument comes from Elkin’s rather extensive description of the easternmost (or following [^15] In the era of visual observation it was common that an object to the west of a point of reference be called a preceding object, one to the east of it a following object, apparently because when the telescope clock was stopped, the Earth’s rotation moved the star field westward.) condensation. Elkin conveys that “till the end of October [this fragment] was a fair object for observation.” However, by November 11 it became “so covered with haze, that it was obvious that soon nothing but the centre of the great mass would be available for a reference point …, and after Nov. 18 … the place of [the following condensation] was more guessed at than actually seen.” This is a consummate account of a protracted flare-up of a fragment that was undergoing a process of continuous loss of mass and should have problems to survive to the next perihelion. Certainly not the behavior of the principal nucleus B!
@@
Ten days later, the appearance of the comet’s chain of fragments did not look radically different (Figure 11). The only easily perceived change was a significant elongation of condensation A, which could be explained by the presence of microscopic dust in its copious ejecta, subjected to high solar radiation pressure accelerations. If so, the center of mass of this condensation should be near its eastern boundary, and the fainter condensation several seconds of arc to the west of it would be in very good agreement with the position of the principal fragment B, derived from Kreutz’s orbital elements. And while there are no signs of fragments C or D, at least one additional condensation is apparent near the western end of the chain, at a location where one would expect fragment E. In the ten days, the cloud of enveloping debris increased its extent from 68′′ to 78′′ and was still more than twice as long as the computed distance between the fragments A and D.
@@ 9. Concluding Remarks
In the other three drawings fewer than four condensations are apparent, most of them marginally, with the principal fragment B enshrouded in the cloud of debris to the extent that it is completely invisible (with one possible exception). Two drawings are dominated by fragment A, which during October was evidently undergoing a protracted flare-up. By the end of the month it was already fading and nearly gone by November 21.

### v2

@@ 8. Fragmented Nucleus of the Comet of 1138
This conclusion is supported by Kreutz’s own orbital computations, which for the three separation distances give, respectively, ${}^{\prime\prime}\!$.8, ${}^{\prime\prime}\!$.2, and ${}^{\prime\prime}\!$.6 (Table 8). The distances BA and BC agree with Elkin’s measurements to $\pm$0′′.5 and only CD is off by nearly 2′′, possibly because of a contribution to D from E. Another strong argument comes from Elkin’s rather extensive description of the easternmost (or following [^15] In the era of visual observation it was common that an object to the west of a point of reference be called a preceding object, one to the east of it a following object, apparently because when the telescope clock was stopped, the Earth’s rotation moved the star field westward.) condensation. Elkin conveys that “till the end of October [this fragment] was a fair object for observation.” However, by November 11 it became “so covered with haze, that it was obvious that soon nothing but the centre of the great mass would be available for a reference point …, and after Nov. 18 … the place of [the following condensation] was more guessed at than actually seen.” This is a consummate account of a protracted flare-up of a fragment that was undergoing a process of relentless loss of mass, with survival problems on the horizon. Certainly not the prospects one would expect for the principal nucleus B!
@@
Ten days later, the appearance of the comet’s chain of fragments did not look radically different (Figure 11). The only easily perceived change was a significant elongation of condensation A, which could be explained by the presence of microscopic dust in its copious ejecta, subjected to high solar radiation pressure accelerations. If so, the center of mass of this condensation should be near its eastern boundary, and the fainter condensation several seconds of arc to the west of it would be in very good agreement with the position of the principal fragment B, derived from Kreutz’s orbital elements. And while there are no signs of fragments C or D, at least one additional condensation is apparent near the western end of the chain, at a location where one would expect fragment E. In the ten days, the cloud of enshrouding debris increased its extent from 68′′ to 78′′ and was still more than twice as long as the computed distance between the fragments A and D.
@@
Figure 13.— Path of the Great Comet of 1106 across the sky (equinox J2000). The positions 1–15 cover the period of time from February 6 through April 17 at a 5-day step, each identified in Table 9. The schematically drawn tail only serves to show the antisolar direction. The path, describing the comet’s sky journey from Aquarius to the belt of Orion, was in its entirety seen from much of the northern hemisphere.
9. Path Across the Sky and Ephemeris
As the representatives of the two foremost populations, the Great Comet of 1106 and the Chinese comet of 1138 (Ho 403) were on a par in terms of their significance in the hierarchy of the Kreutz sungrazer system. Yet, because of the difference in their timing on a scale of less than one thousandth of their orbital period, to the terrestrial observer the first presented a spectacle, the other a run-of-the-mill “broom star.”
Figure 14.— Path of the Chinese comet of 1138 (Ho 403) across the sky (equinox J2000). The positions 1–7 cover the period of time from August 7 through October 6 at a 10-day step (Table 9). The tail is drawn schematically as in Figure 13. The comet’s path begins near Regulus ($\alpha$ Leo) and continues southwards through the constellation Hydra, along the boundary with Sextans.
What were then the paths of the two sungrazers in the sky? Based on the orbits in Table 1, this question is answered in Figures 13 and 14. The ephemerides of the two comets, with the predicted apparent magnitudes derived from the light-curve laws in (4), are in Table 9. The Earth was crossing the comets’ orbital planes on 1106 March 7.9 and 1138 August 20.1 TT, respectively.
@@ 10. Concluding Remarks
In the other three drawings fewer than four condensations are apparent, most of them marginally, with the principal fragment B obscured by the cloud of debris to the extent that it is completely invisible (with one possible exception). Two drawings are dominated by fragment A, which during October was evidently undergoing a protracted flare-up. By the end of the month it was already fading and nearly gone by November 21.